Use the following table to answer the question
Population of Europe by Age Groups
Population (in thousands) 1950 1970 1995
Age 0-14 143,175 166,367 139,464
Age 15-64 359,162 421,432 487,110
Age 65+ 44,981 68,642 101,338
Age 75+ 14,553 22,762 38,139
Total 547,318 656,441 727,912
What was the percentage increase in the age 75+ population of Europe from 1970 to 1995?
A). 65.1%
B).67.6%
C). 63.4%
D). 68.9%

Answer by rajagopalan(174)

The computation of answer is like this:
Base year = 1975
population of Europe in 1995 = 38,139
population of Europe in 1970 = 22,762
Increase in population =38139-22762=15,377
% increase = (15,377/22762)* 100=0.67555*100=67.55
The correct option is 67.6% = (B)
Okay!
Rajagopalan..